OpenStudy (anonymous):

Find the product of z1 and z2 where z1 = 8(cos 40° + i sin 40°), and z2 = 4(cos 135° + i sin 135°).

OpenStudy (anonymous):

\[e^{i \theta}=\cos(\theta)+i \sin(\theta)\]

OpenStudy (anonymous):

???

jimthompson5910 (jim_thompson5910):

if z1 = r1*[cos(theta1) + i*sin(theta1)] and z2 = r2*[cos(theta2) + i*sin(theta2)] then z1*z2 = (r1*r2)*[cos(theta1+theta2) + i*sin(theta1+theta2)]

OpenStudy (anonymous):

Can you send a drawing?

OpenStudy (anonymous):

Use the draw tool.

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

if \[\Large z_1 = r_1[\cos(\theta_1)+i*\sin(\theta_1)]\] and \[\Large z_2 = r_2[\cos(\theta_2)+i*\sin(\theta_2)]\]

jimthompson5910 (jim_thompson5910):

then \[\Large z_1*z_2 = (r_1*r_2)[\cos(\theta_1+\theta_2)+i*\sin(\theta_1+\theta_2)]\]

OpenStudy (anonymous):

Oh, okay.

OpenStudy (anonymous):

What would it look like with the numbers plugged in?

jimthompson5910 (jim_thompson5910):

in this case, r1 = 8, r2 = 4 theta1 = 40, theta2 = 135

OpenStudy (anonymous):

Thanks.